# VeriRAG: A Retrieval-Augmented Framework for Automated RTL Testability Repair

### v1

@@ Abstract
Index Terms:
Design for testability, computer-aided design, large language models, retrieval-augmented generation
@@ I Introduction
Recent advancements in large language models (LLMs) have demonstrated significant potential in enhancing computer-aided design (CAD) workflows, particularly in automated debugging and verification of hardware description languages (HDLs) [1, 2]. However, despite the growing interest in LLM-empowered CAD tools, the domain of design for testability (DFT) remains comparatively underexplored.
@@
Given a Verilog code snippet potentially requiring DFT-compliance code revisions, VeriRAG analyzes the hardware structure of the RTL design and retrieves the most similar example from the VeriDFT dataset based on both hardware structure and DFT errors. The retrieved example, along with our manually edited solution for DFT compliance, serves as a tailored in-context reference to guide the LLM’s correction of the given Verilog code snippet.
Importantly, VeriRAG’s error correction mechanism iteratively refines the RTL design by leveraging compiler diagnostics and DFT error reports generated by electronic design automation (EDA) tools. This iterative process ensures that the modifications effectively address DFT concerns while preserving the design’s synthesizability and logical integrity.
@@
While the above results highlight the potential of VeriRAG, further investigations might be conducted to boost the success rate. First, expanding the VeriDFT dataset to include a broader range of DFT error scenarios and RTL design patterns could provide the LLM with more diverse in-context references, thereby improving its ability to generalize to unseen cases. Second, exploring advanced retrieval techniques, such as graph-based methods to capture deeper structural similarities in hardware designs, could further refine the selection of in-context examples. Third, integrating domain-specific pretraining for LLMs on hardware description languages, particularly with a focus on DFT-related tasks, may significantly boost performance by aligning model representations more closely with this specialized domain. Finally, post-training treatments, such as reinforcement learning with automated EDA grading based on fault coverage metrics, could guide the language model for better results.
@@ II-A LLM for CAD
LLMs are increasingly being used in CAD, supporting tasks ranging from high-level synthesis to RTL bug detection. Recent studies have shown that LLMs can assist with targeted design improvements, automate documentation, and even interface with EDA tools for power and timing analysis based on user inputs [8, 9, 10, 11, 12, 13]. However, most of these efforts have focused on functional correctness, such as verifying data paths or refining structural descriptions, while largely neglecting the critical aspect of testability. Our work addresses this gap by specifically targeting DFT compliance in RTL designs, broadening the scope of LLM-assisted CAD to include comprehensive testability objectives.
@@ II-B DFT at RTL level
Conventional DFT techniques — such as scan-chain insertion [14], boundary scan [15], and built-in self-test [16] — primarily operate at the gate level or during physical implementation. While these methods automate many aspects of test generation and reduce manual effort in late design stages, they do not address defects introduced at the RTL stage. Consequently, designers must rely on linting and static analysis tools to identify code-quality issues. The systematic identification and correction of DFT-related errors, particularly those involving internal clocking or asynchronous resets, remains labor-intensive and requires significant expertise [17].
To address these challenges, we propose an LLM-assisted automated framework enabling the detection and correction of critical DFT issues at the RTL level in a fully automatic manner. Our solution enhances testability without compromising design integrity, substantially reducing the need for manual intervention in DFT debugging. In this work, we focus on four representative DFT error types: asynchronous set/reset inaccessible from primary inputs (ACNCPI), internally generated clock domains (CLKNPI), clock used as data (CDFDAT), and flip-flop-driven clocks (FFCKNP).
@@ II-C RAG for RTL design
RAG combines generative modeling with the retrieval of relevant external knowledge to enhance performance. When employing RAG for our purpose, code templates and bug-fix repositories can be leveraged to guide language model outputs, improving both the precision and correctness of LLM-assisted code repairs [18]. For RTL design, previous retrieval-based methods [19] have shown that proactively identifying and reusing well-validated IP and design blocks can significantly improve development efficiency. However, applying RAG to DFT repair introduces a unique challenge: defining and measuring “similarity” for the precise retrieval of the most relevant reference from the knowledge base. To address this, we propose an autoencoder, trained using the contrastive learning method, to quantify DFT-related hardware similarity for effective retrieval within the RAG framework.
@@ II-D Deep Learning for Hardware and DFT Error Analysis
Deep learning has recently become a powerful tool for hardware analysis, achieving notable results in power modeling [20], defect localization [21], and timing optimization [22, 23]. Most existing methods cast hardware analysis as classification or regression tasks, such as anomaly detection, layout optimization, or performance prediction. This paper focuses specifically on the measurement of similarity among RTL designs with DFT defects. We train a supervised autoencoder using contrastive learning to map Verilog files into a latent embedding space that captures their hardware structures and DFT error types. The similarity of two RTL designs is quantified via the cosine similarity between their embedding vectors.
@@ IV-A Training Autoencoder Network
where $m$ is a hyperparameter representing the fixed margin, describing the targeted minimum distance between two normalized embedding pairs with different labels. $N$ denotes the total number of unique unordered sample-pairs in the batch, i.e. $N=B(B-1)/2$.
@@ VI Conclusions and Outlook
We present VeriRAG, a retrieval-augmented framework designed to tackle DFT error repair at the RTL stage. This framework employs an autoencoder network trained by contrastive learning to identify and retrieve structurally similar reference code from a newly developed Verilog DFT dataset, referred to as VeriDFT. Using the retrieved reference RTL designs, an LLM performs DFT error repairs through an iterative Verilog code revision pipeline. This iterative code-revision process, guided by compiler diagnostics, ensures accurate and reliable error correction.
On a benchmark of 317 tested RTL designs, VeriRAG improves the success rate of DFT repairs by different state-of-the-art LLM models: e.g., from 3% to 40.26% with Grok-3, and from 6.96% to 53.76% with GPT-o1 (13.42-fold and 7.2-fold improvements, respectively). Ablation studies validate the effectiveness of our two key technical contributions: (1) the hardware and DFT similarity-based retrieval scheme, and (2) the iterative code revision pipeline guided by compiler feedback.
There are several promising directions for future work. One open problem revealed in our experiments is the gap between the preliminary success rate and the ultimate success rate. This gap indicates that all these tested LLMs are struggling with writing logically equivalent Verilog codes. This is probably due to these LLMs’ inability to understand subtle circuit design details within these RTL files. We see three avenues to close this gap. First, LLMs fine-tuned, or even pre-trained in the ideal case, on Verilog datasets may attain foundational understanding of hardware structure, hierarchy, and test logic from the massive training data. Second, embedding a lightweight formal-equivalence oracle – e.g., a SAT/SEC engine or incremental LEC – into each repair iteration, turning logical discrepancies into explicit feedback for the LLM. Such “formal-in-the-loop” guidance can bridge the reasoning shortfall without incurring prohibitive runtime overhead and substantially narrow the preliminary-to-ultimate performance gap. Third, our present flow vectorizes from Verilog to JSON may introduce information loss in terms of the hardware’s logical design. A learnable vectorizer/de-vectorizer jointly trained within the network training pipeline – augmented with logic-level signals and equivalence-preserving constraints – could better improve downstream retrieval and repair quality.
On the data side, future work could further enrich VeriDFT. We intend to include more types of DFT violations in VeriDFT and introduce more complex RTL designs with multiple types of DFT errors. Furthermore, future research can strengthen the coverage of the reference dataset in terms of number and diversity by adding more manually edited RTL examples, which can refine the granularity of the retrieval process and result in more reliable DFT error repairs.

### v2

@@ I Introduction
Recent advancements in large language models (LLMs) have demonstrated significant potential in enhancing computer-aided design (CAD) workflows, particularly in automated debugging and verification of hardware description languages (HDLs) [7, 11]. However, despite the growing interest in LLM-empowered CAD tools, the domain of design for testability (DFT) remains comparatively underexplored. Unlike functional verification, which establishes behavioral correctness, DFT at RTL aims to secure observability/controllability and test access properties (e.g., scan readiness, safe clock/reset handling) that enable robust ATPG and manufacturing test.
@@
Given a Verilog code snippet potentially requiring DFT-compliance code revisions, VeriRAG analyzes the hardware structure of the RTL design and retrieves the most similar example from the VeriDFT dataset based on both hardware structure and DFT errors. The retrieved example, along with our manually edited solution for DFT compliance, serves as a tailored in-context reference to guide the LLM’s correction of the given Verilog code snippet. Each reference in the library is paired with a validated fix, which mitigates the risk of misleading edits from mismatched or unverified exemplars.
Importantly, VeriRAG’s error correction mechanism iteratively refines the RTL design by leveraging compiler diagnostics and DFT error reports generated by electronic design automation (EDA) tools. This iterative process ensures that the modifications effectively address DFT concerns while preserving the design’s synthesizability and logical integrity. A repair is accepted only if (i) the design remains synthesizable, (ii) no DFT violations are reported, and (iii) a final logic equivalence check (LEC) passes against the original.
@@ II-A LLM for CAD
LLMs are increasingly being used in CAD, supporting tasks ranging from high-level synthesis to RTL bug detection. Recent studies have shown that LLMs can assist with targeted design improvements, automate documentation, and even interface with EDA tools for power and timing analysis based on user inputs [6, 17, 18, 32, 10, 28]. However, most of these efforts have focused on functional correctness, such as verifying data paths or refining structural descriptions, while largely neglecting the critical aspect of testability [33]. Our work addresses this gap by targeting DFT compliance at the RTL level, where success requires not only compilable edits but also the elimination of DFT violations and preservation of logical equivalence. In contrast to repository- or function-oriented frameworks (e.g., AutoVCoder; RTLRepoCoder) that primarily pursue functional completion or synthesis-aligned correctness, our setting is testability-driven and closes the loop with tool-based DFT diagnostics and a final logic equivalence check (LEC).
@@ II-B DFT at RTL level
Conventional DFT techniques — such as scan-chain insertion [30], boundary scan [4], and built-in self-test [20] — primarily operate at the gate level or during physical implementation. While these methods automate many aspects of test generation and reduce manual effort in late design stages, they do not address defects introduced at the RTL stage. Consequently, designers must rely on linting and static analysis tools to identify code-quality issues. The systematic identification and correction of DFT-related errors, particularly those involving internal clocking or asynchronous resets, remains labor-intensive and requires significant expertise [13].From a process perspective, gate-level DFT insertion cannot reliably repair RTL-origin violations (e.g., internally generated clocks or uncontrollable resets); therefore, catching and fixing such patterns early at RTL is crucial for avoiding repeated synthesis–ATPG iterations and schedule risks. We focus on four representative RTL-level DFT patterns: ACNCPI (asynchronous set/reset not controllable from primary inputs), CLKNPI (internally generated or gated clocks), CDFDAT (a clock net used as data), and FFCKNP (flip-flop–driven clocks).
@@ II-C RAG for RTL design
RAG combines generative modeling with the retrieval of relevant external knowledge to enhance performance. When employing RAG for our purpose, code templates and bug-fix repositories can be leveraged to guide language model outputs, improving both the precision and correctness of LLM-assisted code repairs [16, 22]. For RTL design, previous retrieval-based methods [15] have shown that proactively identifying and reusing well-validated IP and design blocks can significantly improve development efficiency. However, applying RAG to DFT repair introduces a unique challenge: defining and measuring “similarity” for the precise retrieval of the most relevant reference from the knowledge base. In particular, DFT cues (clocks, resets, scan enables) are sparse and topology-dependent, so text-level similarity is insufficient. We therefore adopt a structure-aware retrieval formulation that quantifies hardware similarity with respect to DFT-relevant connectivity, enabling precise selection of validated reference–fix pairs within the RAG loop.
@@ II-D Deep Learning for Hardware and DFT Error Analysis
Deep learning has recently become a powerful tool for hardware analysis, achieving notable results in power modeling [5], defect localization [31], and timing optimization [12, 24]. Most existing methods cast hardware analysis as classification or regression tasks, such as anomaly detection, layout optimization, or performance prediction. In the context of DFT, a complementary line of work learns embeddings of RTL structures so that designs sharing DFT-relevant topologies are close in a latent space; similarity can then be computed (e.g., via cosine distance) to retrieve a small number of validated references that guide subsequent edits. Our study situates DFT-oriented similarity learning within this landscape and employs it strictly as a retrieval component, while the actual code repair is governed by compiler/DFT diagnostics and LEC-based acceptance.
@@ IV-A Training Autoencoder Network
where $m$ is a hyperparameter representing the fixed margin, describing the targeted minimum distance between two normalized embedding pairs with different labels. $N$ denotes the total number of unique unordered sample-pairs in the batch, i.e. N=B(B1)/2N=B(B-1)/2.
@@ VI Conclusions and Outlook
We present VeriRAG, a retrieval-augmented framework for repairing DFT errors at the RTL stage. Concretely, the approach employs an autoencoder trained with contrastive learning to retrieve structurally similar references from a newly developed Verilog DFT dataset, VeriDFT. Leveraging these targeted references, an LLM performs repairs through an iterative Verilog code-revision loop guided by compiler and DFT diagnostics. A candidate fix is accepted only when the design remains synthesizable, contains no DFT violations, and passes a final logic equivalence check (LEC), ensuring that improvements in testability do not compromise functional intent.
[…]
This study has practical boundaries. The dataset focuses on single-violation cases across four representative DFT types, reflecting a controlled setting for isolating effects. Blank-module substitution is applied only when interfaces and connectivity are preserved; cases that would suppress or fabricate diagnostics are excluded, though residual bias cannot be completely ruled out. The reference library is intentionally modest in size to stress retrieval quality and to make failure modes more interpretable.
Looking forward, several directions appear promising. First, domain-adaptive pretraining or fine-tuning of LLMs on Verilog and DFT corpora could strengthen sensitivity to logical-equivalence constraints [21]. Second, lightweight formal-in-the-loop checks (SEC/LEC) at each iteration may convert equivalence gaps into actionable feedback, narrowing the distance between preliminary and ultimate success [9]. Third, moving beyond static Verilog-to-JSON vectorization toward learnable (de)vectorizers that preserve logic-level invariants could improve both retrieval fidelity and edit quality. Finally, expanding the dataset to cover additional DFT types, multi-violation designs, and a larger, more diverse set of validated references would better stress-test retrieval at scale and sharpen external validity. Collectively, these extensions position VeriRAG as a practical foundation for integrating LLMs into DFT-aware RTL workflows.